Help needed desperately !
I have a problem with a recently purchased DHOM 532nm (model DHL-100A PSU, and model DHL-W532-1000mw head).
Power starts off nice and high and well over spec when first switched on but degenerates to less than half a watt within minutes.
If I switch off for a few minutes then switch back on the symptoms repeat themselves.
I can not detect any heating what so ever on the base of the head module so I assume that the peltier/s is/are not operating correctly.
Does anyone have a schematic for this laser since I need to know whether the problem is the temp sensor in the head, the peltier itself or the peltier driver.
I am loathe to take the head apart as dirt could possibly get in since I do not have a class A clean room or anything remotely similar so I would prefer to diagnose the fault from the PSU end of things for the time being.
I feel I have probably been sold a pup since I can not imagine how this type of fault could have possibly have developed during transport.
Does anyone have any thoughts on this ???.
Has anyone ever heard of a thermistor getting "bumped-off" or a TEC circuit giving up the ghost when obviously not powered up ?
The laser was switched on only ONCE for less than a minute using a 4.5volt battery as modulation input to check that it was lasing on arrival and then immediately replaced into its packaging before I got to test it properly with my Ophir power meter five weeks later.
I have already checked for any apparently loose wires and found none in either plug or socket on the PSU.
There is a fully populated 8pin socket on the PSU and a plug on the cable connecting the PSU to the head (which is hard wired into the head through sealed case holes). Since inside the PSU these eight wires split into four two pin connectors on the main controller board I assume the following in absence of further info:
Two of these are the fan supply the other six are probably 2 x pump input, 2 x peltier drive and 2x temp feedback suggesting that there might be only one peltier for the entire optics train!
Although ideally I would expect at least two peltiers, (one for pump and the other for the frequency doubler and stretcher crystals) and two temp feedback devices.
However despite the four way split in the PSU it is conceivable that there are common rails which would allow for another peltier drive and another temp feedback signal especially if the fan supply is shared with the pump input (very dubious but still possible). The internal PSU wiring design and the wire thicknesses (four thick and four thin however make all this somewhat unlikely).
Also since this laser is clearly of Chinese origin the design criteria is more likely to favour price and carefully selected crystals over thermal stability and linear modulation characteristics.
I am somewhat up shit creek without a paddle here !
Even the wiring diagram (or just the pin out) for the head to PSU connection would be a great help since this would enable me to monitor the peltier drive voltage/current and the temp feedback from the head so as to help diagnose the exact nature of the problem.
Does anyone know if DHOM lasers frequently suffer this type of fault ??
Who has repaired/stripped/rebuilt one of these babies?
Any info will be gratefully received either as a post or a PM.
Thanks!
Cheers![]()